is x+7y=8 parellel or perpindicular to y=7x+2
Answer by stanbon(75887)

is x+7y=8 parellel or perpindicular to y=7x+2
---
Solve each for "y" so you can see the slopes.
x + 7y = 8
7y = -x + 8
y = (-1/7)x + (8/7)
slope = -1/7
------------------------
y = 7x+2
slope = 7
================
Product of slopes = (-1/7)7 = -1
so the lines are perpendicular.
======================
Cheers,
Stan H.
==============